ABC has the following accounts in the Property, Plant and Equipment (PPE) section
of its balance sheet: Land, Buildings, and Equipment. Each non-current asset account
has a separate accumulated depreciation contra-account except for Land. ABC uses
the historical cost principle to value its fixed assess after acquisition. ABC completed
the following transactions in fiscal year 2017. ABC has a fiscal year end of 31st
December.
3rd January 2017: ABC exchanged old equipment with accumulated depreciation of
€130,000 (cost of acquisition = €272,000) for similar type of new equipment with a
selling price of €350,000. In addition to the exchange of the old equipment with the
new, ABC paid €200,000 in cash to the supplier of the new equipment. The fair
market value of the equipment given up by ABC cannot be estimated reliably.
1st July 2017: ABC sold the Building A. The Building A appeared with a cost of
€1,310,000 and accumulated depreciation of €260,000 in the balance sheet of ABC
for the prior fiscal year 2016. ABC uses the straight-line depreciation method for
estimating the depreciation expense of its buildings. The management of ABC
estimates a useful life of Building A equal to 40 years and a residual value equal to€310,000. ABC received from the purchaser of the building €230,000 cash and a
€710,500 note receivable.
31st October 2017: ABC made significant capital expenditures, acquiring land and a
new building for a total price of €780,000. Based on estimates of an independent
appraisal, the land had a fair value of €442,200 and the building of €361,800.
31st December 2017: ABC recognized depreciation for its fixed assets based on the
following information:
- The management of ABC expects a useful life for its newly purchased equipment
equal to 5 years and a residual value equal to 10% of its acquisition cost. ABC uses
the double-declining-balance method for estimating the depreciation on its equipment.
- The management of ABC expects a useful life for its newly purchased building
equal to 40 years and a residual value equal to 25% of its acquisition cost. ABC uses
the straight-line depreciation method for estimating the depreciation on its buildings.
During the fiscal year 2018, ABC has not purchased or sold any fixed assets (zero netcapital expenditures).
31st December 2018: ABC recognized depreciation for its fixed assets based on the
following information:
- The management of ABC used the same assumptions for the estimation of
depreciation on its equipment as in fiscal year 2017.- The management of ABC changed its assumptions for the estimation of depreciation
on its building compared to fiscal year 2017. The remaining useful life is estimated as
equal to 45 years and the residual value as equal to €50,000. The changes were
effective from the beginning of the fiscal year 2018. ABC continues to use the
straight-line depreciation method for estimating the depreciation on its buildings.
Required:
Based on the above data, please record the journal entries in ABC’s books for fiscal
years 2017 and 2018.
Answer #1
Date Account Title and Explanation Debit Credit
Jan 2017 Equipment - New 3,50,000.00
Accumulated Depreciation 1,30,000.00
Equipment - Old 2,72,000.00
Cash 2,00,000.00
Profit on disposal 8,000.00
Date Account Title and Explanation Debit Credit
Jul 2017 Accumulated Depreciation 2,60,000.00
Cash 2,30,000.00
Notes receivable 7,10,500.00
Loss on Disposal 1,09,500.00
Building - old 1,310,000.00
Date Account Title and Explanation Debit Credit
Oct 2017 Land 4,42,000.00
Building 3,61,800.00
Cash 7,80,000.00
Gain on disposal 23,800.00

Calculation for Depreciation Equipment

Equipment = 3,50,000.00, Useful life = 5 years, Residual value = 35,000

Depreciation using double declining balance method = 100%/ 5 =20%* 2 = 40%

Depreciation 2017 = (3,50,000 -35,000) * 40/100 =1,26,000

Depreciation 2018 = (3,15,000.00 - 1,26,000) * 40/100 = 1,89,000.00 * 40/100 = 75,600

Calculation for Depreciation Building

Building cost - 3,61,800

Useful life - 40 years

residual value = 25% of cost = 90,450

Depreciation = (Cost - scrap value) / 40 = (3,61,800- 90,450) / 40 = 6784.00

To adjust for 3 months starting from Nov 2017 = 6,784.00 *3/12 = 1696.00

2017 Depreciation = 1696.00

Depreciation for 2018 = 361800 - 50000/45 = 6,929 .00

Date Account Title Debit Credit
Dec 2017 Depreciation Equipment 1,26,000.00
Accumulated Depreciation - Equipment 1,26,000.00
Dec 2018 Depreciation Equipment 75,600.00
Accumulated depreciation - Equipment 75,600.00
Dec 2017 Depreciation Building 1,696.00
Accumulated Depreciation - Building 1696.00
Dec 2018 Depreciation Building 6,929.00
Accumulated Depreciation - Building 6,929.00